# AeroGPT: Leveraging Large-Scale Audio Model for Aero-Engine Bearing Fault Diagnosis

### v1

@@ III-B4 Qualitative Analysis
Beyond quantitative metrics, AeroGPT’s interactive diagnostic capabilities were evaluated to assess practical utility in aerospace maintenance scenarios. Fig. 3 demonstrates how AeroGPT leverages its generative foundation to transform fault diagnosis from isolated classification to interactive consultation, enabling multifaceted analysis of bearing defects.
Instead of producing a simple fault label, AeroGPT generates a detailed fault description when initially queried about defect presence. The framework’s output explicitly identifies the fault as “inner ring indentation” with a severity of “150 mm,” providing immediate context for maintenance personnel. When subsequently questioned about severity, the system demonstrates nuanced assessment capabilities, classifying the defect as “moderate” while noting it “should be monitored closely to prevent further deterioration.” This level of granular evaluation is not typically available in traditional classification outputs, which often lack interpretability and actionable insights.
Moreover, AeroGPT exhibits causal reasoning by establishing explicit connections between vibration patterns and their mechanical implications. When asked about performance impacts, it articulates how inner ring indentation directly causes “increased vibration” and reduced “efficiency and lifespan” of the bearing component. This mechanical insight contrasts sharply with traditional classifiers that provide no explanation of fault mechanisms or consequences, merely indicating fault presence and category.
@@
The framework further demonstrates domain expertise through contextually appropriate technical recommendations. When queried about inspection methodologies, AeroGPT recommends specific diagnostic tools, “a borescope for visual inspection and an ultrasonic device to measure the depth of the indentation”, which are tailored to the identified fault type. Similarly, when asked about factors that might exacerbate the condition, the system identifies “high-speed rotation, heavy loads, or high temperatures” as operational parameters that could accelerate deterioration, providing valuable risk management insights for maintenance scheduling.
AeroGPT’s ability to generate contextually appropriate maintenance recommendations calibrated to aerospace standards represents perhaps its most significant operational advantage. The response prescribing a “maintenance check to inspect the bearing and potentially replace the inner ring” aligns with established protocols for aerospace bearing defects of this category and severity. Furthermore, AeroGPT can incorporate domain-specific contextual information from technical maintenance manuals as additional input, enabling fine-grained recommendation calibration that precisely adheres to manufacturer specifications or organizational maintenance procedures, further enhancing recommendation precision. This multifaceted capability effectively bridges the critical gap between fault detection and maintenance action that typically requires expert human interpretation of classification outputs.
These examples confirm that AeroGPT’s generative approach fundamentally transforms fault diagnosis methodology through a comprehensive set of capabilities spanning detection, assessment, causal analysis, technical guidance, risk evaluation, and intervention planning—all delivered through natural language interaction. The system’s ability to reason about mechanical implications and prescribe appropriate interventions represents a paradigm shift in the interaction between maintenance personnel and diagnostic systems in safety-critical aerospace applications. By providing this complete diagnostic workflow within a conversational interface, AeroGPT significantly reduces the expertise barrier for maintenance personnel, potentially accelerating response times and improving maintenance outcomes in aerospace operations.
III-C Case 2: HIT Bearing Dataset
III-C1 Dataset Descrption
@@
The second dataset utilized was obtained by Hou et al. from the Harbin Institute of Technology (HIT) [35], featuring inter-shaft bearing data from an aero-engine system. Their experimental setup comprises a modified aero-engine with the critical dual-rotor structure preserved. The test rig follows a similar structure with that in the first case, with the low-pressure rotor connected to the high-pressure rotor through a shaft and the inter-shaft bearings subjecting to various fault conditions. The system is equipped with two independent motors for the high and low pressure rotors, allowing for precise control over operational parameters. All signals were sampled at 25kHz, with tests conducted across 28 distinct operating conditions varying both speed (ranging from 1000 r/min to 6000 r/min) and speed ratio between the rotors (1.2 to 1.8). The dataset encompasses 2412 sets of vibration data labeled into three categories (healthy, inner ring defect, and outer ring defect), segmented into 20480-point sequences. The obtained training set consists of 1929 samples and the test set contains 483 samples.
III-C2 Comparison with Existing Methods
@@
TABLE III: Performance comparison of AeroGPT against traditional deep learning models on the HIT dataset.
@@
AeroGPT 100.00% 100.00% 100.00%
@@ III-C3 Qualitative Analysis
The comparative qualitative evaluation illustrated in Fig. 4 demonstrates AeroGPT’s distinctive advantages over both general-purpose large language models and traditional deep learning approaches when diagnosing aero-engine bearing faults.
When presented with identical vibration signals, AeroGPT produces a definitive diagnosis with high specificity, stating unequivocally “Yes, there is an indentation on the inner ring.” This precise identification contrasts sharply with the responses from other large language models, which exhibit significant diagnostic uncertainty despite similar architectural capabilities. The large foundation audio model, Qwen2-Audio-Instruct, produces the tentative assessment “Yes, the signal indicates that there may be a defect,” while the general-purpose model ChatGPT similarly suggests ‘Yes, the yellow part indicates that there may be a defect in the bearing.” The hedging language reflects these models’ lack of specialized knowledge alignment with bearing vibration patterns, resulting in diagnostically ambiguous conclusions that offer limited practical utility for maintenance personnel.
The diagnostic specificity gap becomes even more pronounced when users request detailed fault information. When asked to distinguish between inner and outer ring defects, AeroGPT maintains its diagnostic precision, confidently identifying the inner ring location. In contrast, both Qwen2-Audio and ChatGPT acknowledge a fundamental limitation with the response “There is no sufficient information to determine.” This inability to provide location-specific diagnosis represents a significant practical limitation, as repair protocols and maintenance prioritization depend heavily on defect location within the bearing assembly.
Traditional deep learning methods present an entirely different set of limitations. While potentially achieving comparable classification accuracy, these approaches output numerical probability distributions (e.g., [0.13, 0.82, 0.05]) that require additional post-processing and expert interpretation to translate into actionable diagnoses. The necessity of this translation layer introduces potential inconsistency and significantly limits practical utility in maintenance contexts. The workflow typically involves programming and training a model, obtaining a probability distribution across classes, and then performing post-processing to map the highest probability (0.82 in this case) to a predefined class label (Class 1). Even after this multi-step process, the system remains inherently non-interactive, with the notation “unable to interact with” highlighting a fundamental limitation—these systems cannot engage in follow-up diagnostic discussions or provide explanations for their classifications.
AeroGPT eliminates these limitations through its knowledge-aligned generative approach. By directly outputting precise fault characterizations in natural language and supporting location-specific diagnoses without additional post-processing, it bridges the interpretability gap that typically separates diagnostic systems from maintenance personnel. This capability transforms the diagnostic paradigm from a complex multi-stage analysis requiring specialized expertise to an intuitive conversational process accessible to broader maintenance staff. The directness and specificity of AeroGPT’s diagnostic outputs represent a significant advancement in practical utility for aerospace applications, where maintenance decisions must be made promptly and with high confidence to ensure operational safety.
IV Conclusion
@@
This paper proposed AeroGPT, a novel framework based on large-scale audio model that transfers knowledge from general audio domains to aero-engine bearing fault diagnosis. By recognizing the intrinsic acoustic-like nature of bearing vibration signals, AeroGPT addresses fundamental limitations in current aero-engine fault diagnosis approaches. The framework’s two key innovations, Vibration Signal Alignment (VSA) and Generative Fault Classification (GFC), provide a systematic methodology for adapting general audio knowledge to domain-specific vibration patterns while enabling direct generation of interpretable diagnostic outputs. Through comprehensive experimental validation on two aerospace bearing datasets, AeroGPT achieved exceptional performance with 98.94% accuracy on the DIRG dataset and perfect 100% classification on the HIT bearing dataset, surpassing traditional deep learning approaches. The qualitative analysis further demonstrated AeroGPT’s unique capability to provide definitive, specific fault characterizations in natural language, contrasting sharply with the uncertainty exhibited by general-purpose language models and the post-processing requirements of conventional classification approaches. This transformation of fault diagnosis from complex analysis to intuitive conversation significantly enhances practical utility in aerospace maintenance contexts, where rapid, accurate interpretation is essential for preventing catastrophic failures. By eliminating the need for post-processing and providing interactive, interpretable, and actionable diagnostics, AeroGPT represents a significant advancement in aerospace reliability engineering, highlighting the substantial potential of large-scale audio models to revolutionize fault diagnosis in industrial settings.

### v2

@@ IV-B4 Qualitative Analysis
Beyond quantitative metrics, AeroGPT’s interactive diagnostic capabilities were evaluated to assess practical utility in aerospace maintenance scenarios. Fig. 3 demonstrates how AeroGPT transforms fault diagnosis from isolated classification to interactive consultation.
[…]
@@
The framework further demonstrates domain expertise through contextually appropriate technical recommendations. When queried about inspection methodologies, AeroGPT recommends “a borescope for visual inspection and an ultrasonic device to measure the depth of the indentation.” When asked about exacerbating factors, it identifies “high-speed rotation, heavy loads, or high temperatures” as operational parameters that could accelerate deterioration. Additionally, the response prescribing a “maintenance check to inspect the bearing and potentially replace the inner ring” aligns with established aerospace protocols, effectively bridging the gap between fault detection and maintenance action.
The interactive dialogue capability demonstrates the advantage of the generative formulation: rather than outputting probability distributions requiring expert interpretation, AeroGPT directly produces human-readable diagnoses and engages in follow-up reasoning about fault mechanisms, severity assessment, and maintenance recommendations. This end-to-end interpretability eliminates the need to translate model outputs into maintenance actions. When combining with techniques such as RAG, AeroGPT can access a vast knowledge base to provide more accurate and contextually relevant responses.
IV-C Case 2: HIT Bearing Dataset
IV-C1 Dataset Description
@@
The second dataset was obtained by Hou et al. from the Harbin Institute of Technology (HIT) [12], featuring inter-shaft bearing data from an aero-engine system. Their experimental setup comprises a modified aero-engine with the critical dual-rotor structure preserved. The test rig follows a similar structure to that in the first case, with the low-pressure rotor connected to the high-pressure rotor through a shaft and the inter-shaft bearings subjected to various fault conditions. The system is equipped with two independent motors for the high and low pressure rotors, allowing for precise control over operational parameters. All signals were sampled at 25 kHz, with tests conducted across 28 distinct operating conditions varying both speed (ranging from 1000 rpm to 6000 rpm) and speed ratio between the rotors (1.2 to 1.8). The dataset encompasses 2412 sets of vibration data labeled as three categories (healthy, inner ring defect, and outer ring defect), segmented into 20480-point sequences. The obtained training set consists of 1929 samples and the test set contains 483 samples.
IV-C2 Comparison with Existing Methods
@@
TABLE III: Performance comparison of AeroGPT against both fault diagnosis models and general-purpose deep learning models on the HIT dataset.
@@
Fault Diagnosis Models
AeroGPT (Ours) 100.00% 100.00% 100.00%
LiConvFormer 99.79% 99.80% 99.79%
TFN 98.58% 98.58% 98.58%
DCNDSC 99.39% 99.40% 99.39%
ILDM 98.76% 98.77% 98.76%
@@
MPSO-ACBCNN 99.79% 99.79% 99.79%
TAUN 99.38% 99.39% 99.38%
DSRSN 98.38% 98.39% 98.38%
SPCFormer 98.35% 98.37% 98.34%
MRCNN-LSTM 94.42% 94.48% 94.44%
General Purpose Models
@@ IV-C3 Qualitative Analysis
The comparative qualitative evaluation illustrated in Fig. 4 demonstrates AeroGPT’s distinctive advantages over both general-purpose large language models and traditional deep learning approaches.
When presented with the same vibration signal, AeroGPT produces a definitive diagnosis stating “Yes, there is an indentation on the inner ring,” while general-purpose models can only make tentative assessments. Qwen2-Audio-Instruct produces the tentative “Yes, the signal indicates that there may be a defect,” and ChatGPT suggests “Yes, the yellow part indicates that there may be a defect in the bearing.” When asked to distinguish between inner and outer ring defects, AeroGPT confidently identifies the inner ring location, whereas both Qwen2-Audio and ChatGPT respond “There is no sufficient information to determine.” Traditional deep learning methods present different limitations: they output numerical probability distributions (e.g., [0.13, 0.82, 0.05]) requiring post-processing and expert interpretation, and remain inherently non-interactive.
This comparison highlights the core advantage of vibration-text alignment. While general-purpose audio models possess acoustic understanding capabilities, they lack domain-specific grounding for precise fault localization. The VSA stage bridges this gap by learning vibration-language correspondences, enabling AeroGPT to distinguish between fault locations that general models cannot differentiate. While conventional deep learning methods can also perform well on this task, they require post-processing and expert interpretation. These results validate AeroGPT’s capabilities in achieving both high accuracy and interactivity in industrial fault diagnosis.
V Further Discussion
V-A Computational Cost and Inference Latency
For practical deployment considerations, we report the computational requirements and inference performance of AeroGPT. The foundation model comprises a pre-trained audio encoder with 124 million parameters and a large language model with 7 billion parameters. Using LoRA-based adaptation, only approximately 14 million parameters (0.2% of the total) require training, significantly reducing computational requirements compared to full fine-tuning.
Under our experimental setup (NVIDIA A10 GPU with 24 GB VRAM, batch size 1), AeroGPT achieves the following performance characteristics: the average inference speed is 21.35 tokens per second, with an average time-to-first-token latency of 1.9 seconds, which means that simple fault diagnosis tasks can be completed within 5 seconds. The memory footprint during inference depends on the KV cache size, and a single GPU with 24 GB VRAM is sufficient for deployment. For training, the LoRA adaptation requires approximately 17 GB VRAM with gradient accumulation, enabling fine-tuning on a single GPU rather than requiring multi-GPU clusters.
These characteristics indicate that AeroGPT is suitable for offline diagnostic applications and near-real-time monitoring scenarios where latency requirements are on the order of seconds rather than milliseconds. For applications requiring faster response times, potential optimizations include model quantization, speculative decoding, or distillation to smaller architectures.
V-B Potential Misdiagnosis Risks
For deployment in safety-critical aerospace applications, careful consideration of misdiagnosis risks is essential. Several potential failure modes must be considered. False negatives, i.e., missed faults, could lead to undetected bearing degradation and potential in-service failures; our experimental results show that AeroGPT achieves high recall (99.07% on DIRG defective samples), but any missed detection in safety-critical applications warrants additional safeguards. False positives, i.e., false alarms, could trigger unnecessary maintenance actions, increasing operational costs; the high precision (99.26% on DIRG defective samples) mitigates but does not eliminate this risk. Out-of-distribution (OOD) conditions, including novel fault types, operating regimes not seen during training, or sensor configurations different from training data, may produce unreliable predictions. Additionally, sensor faults such as corrupted, noisy, or missing input signals could lead to erroneous diagnoses.
V-C Safe Integration into Existing Workflows
To safely and responsibly make use of AeroGPT in safety-critical aerospace applications, it is crucial to consider the potential risks and liabilities associated with its deployment. We emphasize that AeroGPT should be deployed as a decision-support module within an engine health monitoring workflow, not in the closed-loop engine control path. The framework is designed to augment rather than replace expert judgment in maintenance decision-making.
To mitigate misdiagnosis risks, we recommend several strategies for safe integration. First, an out-of-distribution (OOD) detection module should be placed upstream of the AeroGPT framework to filter out input signals that deviate significantly from the training distribution, such as those caused by sensor malfunctions, extreme operating conditions, or unknown interference, ensuring that the model is only queried with valid, high-fidelity data within its operational design domain. Second, a lightweight deterministic deep learning model can be deployed in parallel with AeroGPT as a cross-verification agent. If the generative output conflicts with the classification result of the deterministic model, the system should flag the discrepancy as a high-risk event for immediate investigation. Third, rule-based consistency checks against physics-based models, historical maintenance records, or redundant sensor readings can validate AeroGPT’s outputs. Fourth, a human-in-the-loop review process should always be maintained for all critical maintenance decisions, with AeroGPT serving as a first-pass screening tool that prioritizes cases for expert attention. Fifth, online learning or periodic model updates should be implemented to adapt AeroGPT to emerging distribution shifts and novel fault patterns.
By positioning AeroGPT as an assistive diagnostic tool within a comprehensive health monitoring workflow that includes multiple redundant checks and human oversight, the framework can provide significant value in accelerating fault identification and reducing diagnostic workload while maintaining the safety standards required for aerospace applications.
VI Conclusion and Future Work
@@
This paper proposes AeroGPT, a novel framework based on a large-scale audio model that transfers knowledge from the general audio domain to aero-engine bearing fault diagnosis. By recognizing the intrinsic acoustic-like nature of bearing vibration signals, AeroGPT addresses fundamental limitations in current aero-engine fault diagnosis approaches. The framework’s two key innovations, Vibration Signal Alignment (VSA) and Generative Fault Classification (GFC), provide a systematic methodology for adapting general audio knowledge to domain-specific vibration patterns while enabling direct generation of interpretable diagnostic outputs. Through comprehensive experimental validation on two aerospace bearing datasets, AeroGPT achieved exceptional performance with 98.94% accuracy on the DIRG dataset and perfect classification accuracy on the HIT bearing dataset, surpassing traditional deep learning approaches. The qualitative analysis further demonstrated AeroGPT’s unique capability to provide definitive, specific fault characterizations in natural language, contrasting sharply with the uncertainty exhibited by general-purpose language models and the post-processing requirements of conventional classification approaches. This transformation of fault diagnosis from complex analysis to intuitive conversation significantly enhances practical utility in aerospace maintenance contexts, where rapid, accurate interpretation is essential for preventing catastrophic failures. By eliminating the need for post-processing and providing interactive, interpretable, and actionable diagnostics, AeroGPT represents a significant advancement in aerospace reliability engineering, highlighting the substantial potential of large-scale audio models to revolutionize fault diagnosis in industrial settings. Furthermore, we provide a detailed discussion on practical deployment considerations, including computational cost and inference latency, potential misdiagnosis risks, and strategies for safe integration into existing engine health monitoring workflows.
The promising performance of AeroGPT establishes a strong foundation for several exciting avenues of future research. A key direction for advancement is to broaden the scope of its operational validation. While the current study demonstrates efficacy on two aerospace engine datasets, future work could aim to assess the model’s performance under a more extensive range of operating conditions, including extreme speeds, heavy loads, and transient states. This will provide a more comprehensive understanding of its reliability in real-world scenarios. Another important area for enhancement involves strengthening the model’s generalization capabilities across diverse hardware and sensing configurations. Future investigations can focus on the model’s adaptability to variations in bearing types, sizes, and manufacturers, as well as different sensor placements and data acquisition systems, which can introduce significant data distribution shifts.